OpenStudy (anonymous):

If the first step in the solution of the equation -8 - 7x = -5x - 10 is "add 10," then what should the next step be?

jhonyy9 (jhonyy9):

-8-7x = -5x-10 add 10 -8+10-7x = -5x-10+10 2-7x = -5x the next step will be add 7x so what will get ?